OpenStudy (anonymous):

Fayed tipped the pizza delivery driver for pizzas that cost . Approximately what percent did he tip the driver? Round your answer to the nearest percent.

OpenStudy (anonymous):

Fayed tipped the pizza delivery driver $2 for pizzas that cost $18 . Approximately what percent did he tip the driver? Round your answer to the nearest percent.

OpenStudy (anonymous):

Divide 2 / 18

OpenStudy (anonymous):

Then multiply by 100.

OpenStudy (anonymous):

so 900

OpenStudy (anonymous):

No..you divided 18 / 2, we're dividing 2 / 18

OpenStudy (anonymous):

should it be \[=\frac{ 2 }{ 18+2 }\times100\]

OpenStudy (anonymous):

Oh..

OpenStudy (anonymous):

so 11.11 repeating

OpenStudy (anonymous):

So you're actually dividing 2 / 20

OpenStudy (anonymous):

10%

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (anonymous):

2 / 20 = 0.10 0.10 * 100 = 10%

OpenStudy (anonymous):

what if it isn't one of my answers?

OpenStudy (anonymous):

A. 2% B. 9% C. 11% D. 16%

OpenStudy (anonymous):

Hmm..maybe I was correct in the first place.. 11%

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i wasn't sure about \[\frac{ 2 }{ 18+2 }\times100\]

OpenStudy (anonymous):

Thnx @iGreen.

OpenStudy (anonymous):

Np